OpenStudy (anonymous):

Write the equation of a circle with center at (-3, -4) and a radius of 3. A. (x - 3)2 + (y - 4)2 = 9 B. (x + 3)2 + (y + 4)2 = 9 C. (x + 3)2 + (y + 4)2 = 3 D. (x - 3)2 + (y - 4)2 = 3

OpenStudy (johnweldon1993):

Equation of a circle \[(x - h)^2 + (y - k)^2 = r^2\] Center = (h,k) so you replace the 'h' with -3 and the 'k' with -4 r = radius....so you replace 'r' with 3 so which does this look like?

OpenStudy (anonymous):

is it A

OpenStudy (anonymous):

NO D